# Towards An Analytical Framework for Dynamic Potential Games

### v2

@@ Appendix A Proofs of Technical Lemmas
A.2 Proofs of Lemmas 5.1, 5.2 and 5.3
Proof of Lemma 5.1.
@@
For Item 1, observe that by Lemma 5.1 and (4.8), for any given $\phi^{\prime\prime}_{j}\in\operatorname*{span}(\pi_{j})$, the map $\phi^{\prime}_{i}\mapsto\big{(}F_{B},F_{\Sigma}\big{)}\left(s,X^{t,x,\phi}_{s},\frac{\delta X^{t,x}}{\delta\phi_{i}}(\phi;\phi^{\prime}_{i}),\frac{\delta X^{t,x}}{\delta\phi_{j}}(\phi;\phi^{\prime\prime}_{j}),\phi^{\prime}_{i},\phi^{\prime\prime}_{j}\right)$ is linear. This along with the linearity of the SDE (4.7) shows that $\phi^{\prime}_{i}\mapsto\frac{\delta^{2}X^{t,x}}{\delta\phi_{i}\delta\phi_{j}}(\phi;\phi^{\prime}_{i},\phi^{\prime\prime}_{j})$ is linear. Similar arguments yield the linearity of $\phi^{\prime\prime}_{j}\mapsto\frac{\delta^{2}X^{t,x}}{\delta\phi_{i}\delta\phi_{j}}(\phi;\phi^{\prime}_{i},\phi^{\prime\prime}_{j})$.
@@
Finally, for all $\varepsilon\in[0,1)$, let $\phi^{\varepsilon}=(\phi_{j}+\varepsilon(\phi^{\prime\prime}_{j}-\phi_{j}),\phi_{-j})$. By Lemma 5.1, $\frac{\delta X^{t,x}}{\delta\phi_{i}}(\phi^{\varepsilon};\phi^{\prime}_{i})$ satisfies
@@
Proof of Lemma 5.3.
For Item 1, the linearity of $\frac{\delta\alpha^{t,x}}{\delta\phi_{i}}(\phi;\cdot)$ follows from (4.9) and the linearity of $\frac{\delta X^{t,x}}{\delta\phi_{i}}(\phi;\cdot)$ (see Lemma 5.1). For each $\varepsilon\in[0,1)$, let $\phi^{\varepsilon}=(\phi_{i}+\varepsilon(\phi^{\prime}_{i}-\phi_{i}),\phi_{-i})$ and observe that
@@
For Item 2, the bilinearity and symmetry of $\frac{\delta^{2}\alpha^{t,x}}{\delta\phi_{i}\delta\phi_{j}}(\phi;\cdot,\cdot)$ follow directly from (4.10) and Lemmas 5.1 and 5.2. For each $\varepsilon\in[0,1)$, let $\phi^{\varepsilon}=(\phi_{j}+\varepsilon(\phi^{\prime}_{j}-\phi_{j}),\phi_{-j})$ and
$\frac{\delta\alpha^{t,x}}{\delta\phi_{i}}(\phi^{\varepsilon};\phi^{\prime}_{i})_{s}$ $=(\partial_{x}\phi^{\varepsilon})(s,X^{t,x,\phi^{\varepsilon}}_{s})\frac{\delta X^{t,x}}{\delta\phi_{i}}(\phi^{\varepsilon};\phi^{\prime}_{i})_{s}+E_{i}\phi^{\prime}_{i}(s,X^{t,x,\phi^{\varepsilon}}_{s}),$ (A.1)

### v3

@@ Appendix B Proofs of Technical Lemmas
B.2 Proofs of Lemmas A.1, A.2 and A.3
Proof of Lemma A.1.
@@
For Item 1, observe that by Lemma A.1 and (4.8), for any given $\phi^{\prime\prime}_{j}\in\operatorname*{span}(\pi_{j})$, the map $\phi^{\prime}_{i}\mapsto\big{(}F_{B},F_{\Sigma}\big{)}\left(s,X^{t,x,\phi}_{s},\frac{\delta X^{t,x}}{\delta\phi_{i}}(\phi;\phi^{\prime}_{i}),\frac{\delta X^{t,x}}{\delta\phi_{j}}(\phi;\phi^{\prime\prime}_{j}),\phi^{\prime}_{i},\phi^{\prime\prime}_{j}\right)$ is linear. This along with the linearity of the SDE (4.7) shows that $\phi^{\prime}_{i}\mapsto\frac{\delta^{2}X^{t,x}}{\delta\phi_{i}\delta\phi_{j}}(\phi;\phi^{\prime}_{i},\phi^{\prime\prime}_{j})$ is linear. Similar arguments yield the linearity of $\phi^{\prime\prime}_{j}\mapsto\frac{\delta^{2}X^{t,x}}{\delta\phi_{i}\delta\phi_{j}}(\phi;\phi^{\prime}_{i},\phi^{\prime\prime}_{j})$.
@@
Finally, for all $\varepsilon\in[0,1)$, let $\phi^{\varepsilon}=(\phi_{j}+\varepsilon(\phi^{\prime\prime}_{j}-\phi_{j}),\phi_{-j})$. By Lemma A.1, $\frac{\delta X^{t,x}}{\delta\phi_{i}}(\phi^{\varepsilon};\phi^{\prime}_{i})$ satisfies
@@
Proof of Lemma A.3.
For Item 1, the linearity of $\frac{\delta\alpha^{t,x}}{\delta\phi_{i}}(\phi;\cdot)$ follows from (4.9) and the linearity of $\frac{\delta X^{t,x}}{\delta\phi_{i}}(\phi;\cdot)$ (see Lemma A.1). For each $\varepsilon\in[0,1)$, let $\phi^{\varepsilon}=(\phi_{i}+\varepsilon(\phi^{\prime}_{i}-\phi_{i}),\phi_{-i})$ and observe that
@@
For Item 2, the bilinearity and symmetry of $\frac{\delta^{2}\alpha^{t,x}}{\delta\phi_{i}\delta\phi_{j}}(\phi;\cdot,\cdot)$ follow directly from (4.10) and Lemmas A.1 and A.2. For each $\varepsilon\in[0,1)$, let $\phi^{\varepsilon}=(\phi_{j}+\varepsilon(\phi^{\prime}_{j}-\phi_{j}),\phi_{-j})$ and
$\frac{\delta\alpha^{t,x}}{\delta\phi_{i}}(\phi^{\varepsilon};\phi^{\prime}_{i})_{s}$ $=(\partial_{x}\phi^{\varepsilon})(s,X^{t,x,\phi^{\varepsilon}}_{s})\frac{\delta X^{t,x}}{\delta\phi_{i}}(\phi^{\varepsilon};\phi^{\prime}_{i})_{s}+E_{i}\phi^{\prime}_{i}(s,X^{t,x,\phi^{\varepsilon}}_{s}),$ (B.1)